# Acceptor and donor impurity levels in hexagonal-diamond silicon

### v1

@@ Abstract
Recent advances in the characterization of hexagonal-diamond silicon (2H-Si) have shown that this material possesses remarkably different structural, electronic and optical properties as compared to the common cubic-diamond (3C) polytype. Interestingly, despite the wide range of physical properties analyzed, to date no study has investigated impurity energy levels in 2H-Si. Here, we present results of ab initio DFT simulations to describe the effect of p- and n-type substitutional doping on the structural and electronic properties of hexagonal-diamond Si (2H-Si). We first provide a detailed analysis on how a given impurity can assume a different local symmetry depending on the host crystal phase. Then, by studying neutral and charged dopants, we carefully estimate donors and acceptors transition energy levels in 2H-Si and compare them with the cubic-diamond (3C) case. In the case of acceptors, the formation energy is always lower in 2H-Si and is associated to a shallower charge transition level with respect to 3C-Si. On the other hand, donors prefer the cubic phase and have transition energies smaller with respect to 2H-Si. Finally, by employing a simple model based on the 2H/3C band offset diagram, we prove the physical validity of our findings and we show how holes can be used to stabilize the 2H-Si phase.
@@ I Introduction
Recent developments and advances in semiconductor growth methods demonstrated that the hexagonal-diamond polytype of Si, Ge and SiGe alloys (2H in the Rasmdell notation Ramsdell (1947)) can be fabricated in the form of nanowires Galvão Tizei and Amato (2020); Fadaly et al. (2020); Vincent et al. (2014); Tang et al. (2017); Hauge et al. (2015). This crystal structure presents a $C_{3v}$ symmetry (P63/mmc space group) with three equal length bonds in the c-plane (0001) and a larger one along the [0001] direction of the hexagonal crystal cell. Compared to the 3C cubic-diamond phase (Fd-3m group symmetry) where four equal length bonds induce a $T_{d}$ symmetry, in the 2H phase the atomic ordering along the $\langle 111\rangle$ cubic direction changes from an ABC to an ABAB stacking.
@@
The main reason for studying these novel group IV polytypes lies in the possibility to observe unexplored electronic and optical properties with respect to those of their 3C counterparts. To give an example, Ge, which in its cubic form is an indirect semiconductor, has a direct band gap in its hexagonal polytype Rödl et al. (2019); Kaewmaraya et al. (2017). Yet, hexagonal SiGe alloys Hauge et al. (2017) and quantum wells Peeters et al. (2024) have recently attracted a great attention because of their direct band gap and the experimentally observed light emission that can be tailored via the Ge concentration and quantum confinement effects Fadaly et al. (2020); Borlido et al. (2021); Peeters et al. (2024). In the case of Si, while the cubic phase has an indirect band gap of 1.12 eV with a conduction band minimum close to the X point, the 2H structure presents an indirect band gap of about 0.95 eV and a conduction band minimum at the L point Rödl et al. (2015); Keller et al. (2023). The direct band gap is also reduced from 3.40 eV to 1.63 eV Rödl et al. (2015). Indeed, 2H-Si presents a larger optical absorption spectra in the visible range than 3C-Si Amato et al. (2016), making it a good candidate for solar panel materials. Furthermore, 3C/2H Si junctions are also of interest for photovoltaics due to their intrinsic tendency to promote charge carrier separation Amato et al. (2016); Kaewmaraya et al. (2017); Belabbes et al. (2022).
@@
It is well known that most technologies based on carrier transport in semiconductors (e.g. electronics, embedded systems, light emitting devices and solar cells) become functional only if the material can be appropriately doped. Impurities play a crucial role as they allow to modulate the carrier concentration around the Fermi level. Furthermore, in different type of semiconductor compounds, doping can be employed to stabilize one polytype with respect to the others Heine et al. (1991); Dalpian et al. (2006); Dalpian and Wei (2004); Liu et al. (2024). For instance, Cai et al. Cai et al. (2021) recently demonstrated that the direct band gap 2H-Ge phase can be stabilized by incorporating electrons into the 3C-Ge crystal. Finally, most of the experimental grown nanowires can contain significant concentration of extrinsic dopants due to the substrate atomic contamination. This is, for instance, the case of 2H-Ge and 2H-SiGe alloys Fadaly et al. (2020) fabricated with the crystal transfer methods which usually present a high concentration of As atoms ( $\sim$ 1018 cm-3).
@@
Although the behaviour of impurities in 3C-Si has been widely discussed, to date no study has investigated impurity energy levels in 2H-Si. The experimental characterization of dopants in this polytype still remains challenging because of the difficulties in its fabrications methods. Ab-initio theoretical approaches can be hence an efficient way to investigate this question. A few studies focused on the analysis of the stability of single-point and neutral substitutional defects in 2H-Si Amato et al. (2019, 2020); Sun et al. (2021) but they were essentially limited to the structural stability of defects. Here, we present a comprehensive study carried out via ab initio DFT simulations on the effect of p- and n-type substitutional defects on the structural and electronic properties of 2H-Si. We first provide a detailed analysis on how a given impurity can assume a different local symmetry depending on the host crystal phase. Then, by studying neutral and charged dopants, we carefully estimate donors and acceptors thermodynamic transition levels in 2H-Si and compare them with the 3C-Si case. Finally, by employing a simple model based on the 2H/3C band offset diagram, we show how the carrier concentration can be used to stabilize one phase with respect to the other.
@@ II Methodology
All simulations were performed by spin-polarized ab initio DFT as implemented in the SIESTA code Soler et al. (2002), using the generalized gradient approximation (GGA) with the Perdew, Burke, Ernzerhof (PBE) functional Perdew et al. (1996). Since the diameter of the 2H nanowires grown experimentally is usually larger than 100 nm, we considered bulk supercells as an accurate and reliable approximation for such systems. Indeed at such large diameters, quantum confinement effect on the band gap can be neglected Amato et al. (2012, 2010); Mohammad (2014). Norm-conserving pseudopotentials of Trouillier-Martins type were chosen to replace the core electrons. Valence electrons were instead represented using a double-$\zeta$ (DZ) polarized basis set for Si, and double-$\zeta$ plus two polarization orbitals (DZP) basis sets for impurities. We considered p-type (B, Al, Ga, In) and n-type (P, Sb, As) as common substitutional dopants in 2H-Si. Lattice parameters and band gap of pristine bulk 3C- and 2H-Si were calculated by sampling the Brillouin zone using a 8$\times$8$\times$8 Monkhorst-Pack k-points grid Monkhorst and Pack (1976). Table 1 presents the converged results of the lattice parameter and the band gap calculations for both the phases. As expected, both quantities are affected by the error associated to the GGA exchange-correlation functional. We used a denser grid of 25$\times$25$\times$25 k-points to calculate the dielectric constant of both phases. This quantity was needed for the charge correction scheme described below Freysoldt et al. (2009). The values of the in-plane and out-of-plane components, reported in Table 1, are in perfect agreement with other DFT calculations Baroni and Resta (1986); Zheng et al. (2017). For the doped systems, a supercell of 512 (a 4$\times$4$\times$4 replica of the 8-atoms conventional supercell with orthogonal cell vectors) and of 500 atoms (a 5$\times$5$\times$5 replica of the primitive cell) were used for the 3C and 2H phases, respectively. This choice has guaranteed the convergence of the formation energy value for both neutral and charged impurities as it eliminated the spurious interaction between defect periodical images. Such doped supercells correspond to a dopant concentration between 1019 cm-3 and 1020 cm-3, a value that is one order of magnitude larger with respect to the impurity concentration usually measured in 2H-samples Fadaly et al. (2020). A 2$\times$2$\times$2 k-grid and a real space grid cut-off of 500 Ry were considered to ensure the proper convergence of the results. Self-consistency field cycles were converged with a tolerance of $10^{-5}$ eV. Atomic positions were optimized until the forces were less than $10^{-2}$ eV/Ang and the stress lower than $10^{-1}$ GPa. Relaxation was performed adopting a conjugate gradient algorithm that calculates DFT forces through the Hellman-Feynman theorem. Charged structures were relaxed starting from the neutral defect optimized structure. A compensating jellium background was added to avoid divergences due to the defect-defect long-range electrostatic interaction between charged periodic cells.
@@
For all the considered impurities, we calculated their formation energy, $E_{form}$, in a given crystal phase (2H or 3C). This quantity is defined as the energy required to replace a host atom with a neutral or charged defect. In the Zhang and Nortrup formalism Zhang and Northrup (1991), it reads as follows:
@@
where $E_{total}^{D}$ is the total energy of the doped system, $E_{total}^{prist}$ is the total energy of the pristine system, $\mu_{0}$ is the chemical potential of Si in a given phase while $\mu_{D}$ is that of the defect. $\mu_{0}$ is here defined as the total ground energy per atom of a pristine Si cell whereas $\mu_{D}$ is calculated as the total energy of the isolated impurity atom. However, it is worth to note that there is no univocal criteria to calculate the chemical potential of an impurity as this quantity strongly depends on the synthesis conditions. As a consequence, the calculated value of the formation energy can assume a wide range of values depending on the way $\mu_{D}$ is defined. Nevertheless, since we are interested in the relative stability of a dopant among the two phases, the chemical potential $\mu_{D}$ erases if we refer to the difference between formation energies, $\Delta E^{2H-3C}_{form}=E_{form}^{2H}-E_{form}^{3C}$ rather than in $E_{form}$ value in itself (this also holds because the difference of $\mu_{0}$ among the two phases can be neglected). Yet, referring back to Eq. (1), the term $q(\epsilon_{VBM}+\mu_{e})$ is the charge multiplied by the electron chemical potential, $\mu_{e}$, using as energy reference the valence band maximum, $\epsilon_{VBM}$. $\mu_{e}$ varies from 0 to the electronic band gap. Finally, the term $E_{corr}$ corresponds to the Freysoldt-Neugebauer-Van de Walle’s finite-size correction Freysoldt et al. (2009) that cancels artificial interactions between charged defects and that has been proven to give very accurate results for different type of defects Freysoldt et al. (2014). As other proposed schemes Dabo et al. (2008); Kumagai and Oba (2014); Freysoldt et al. (2014); Naik and Jain (2018), this approach implies the solution of the Poisson equation for a model system and the alignment of the electrostatic potentials to derive the formation energy correction. In particular, solving the Poisson equation for the periodic model potential needs the calculation of the dielectric tensor profile of the material Naik and Jain (2018), whose static components are presented in Table 1.
@@
Transition levels $\epsilon(q_{1}/q_{2})$ correspond to thermal ionization energies that can be measured by deep level transient spectroscopy (DLTS) Willardson et al. (1998). We employed Eq. (2) to calculate both acceptors, $\epsilon_{A}(q_{1}/q_{2})$ and donor transition levels $\epsilon_{D}(q_{1}/q_{2})$. It is worth to note that transition levels are not affected by the dopant chemical potential which assumes the same value for the charged and neutral impurity and hence cancels out in Eq. (2).
In the following analysis, Eq. (2), which depends on the valence band maximum of each phase taken separately, allows to compare the doping with a given impurity of the 2H and the 3C phase of Si. On the other hand, to give an accurate quantitative estimation of transition levels, in particular in the case of the phase coexistence Vincent et al. (2014); Tang et al. (2017); Amato et al. (2016), we align $\epsilon$($q_{1}$/$q_{2}$) to a common reference level, i.e. the vacuum level $\Phi$ (see SM for the details of the calculations). Indeed, many works have demonstrated that, even if within a semi-local approach, energy levels of dopants generally remain preserved if they are referred to a common well defined potential Alkauskas and Pasquarello (2011); Komsa et al. (2010); Alkauskas et al. (2008). This choice improves the quantitative comparison among the two phases. On the basis of this consideration, Eq. (2) can be modified as:
@@ III Structural properties
We start our discussion by a detailed structural analysis of the relaxed doped structures. Table 2 and 3 report the first-neighbors bond lengths of group III and group V impurities in the cubic and hexagonal crystals. In the 3C-Si, both neutral donors and acceptors maintain the tetrahedral symmetry ($T_{d}$) forming four first-neighbor bonds of equal length. Variations in bond lengths compared with the pristine system are uniform and depend on the mismatch between the covalent radius of the impurity and that of the host element, as already reported by Amato et al. Amato et al. (2019, 2020). The bond contractions and expansions discussed here agree with previous results. As for smaller defects, such as B and P, the first neighbors impurity bonds are compressed in respect to those of the pristine cell (see first block of Table 2 and  3) while for all the other reported defects (Al, Ga, In, Sb and As in Table 2 and  3) they tend to expand since their covalent radius is larger than those of Si. More in general, we can say that impurities in the 3C-Si phase induce an expansion (contraction) of their first neighbors bonds and, consequently, there is a noticeable contraction (expansion) between the first and second neighbors atomic layers that gradually fades within the crystal. The charged impurities also preserve the cubic $T_{d}$ symmetry. As for B-, the extra electron does not induce any appreciable change in the bond distances (see first block of Table 2). Al-, Ga-, and As- first neighbors distances are only slightly reduced compared to the neutral case (see Table 2). These tiny bond length variations are proportional to the impurity size, being more pronounced for larger dopants. Therefore, adding an extra electron to large radius acceptors (Al, Ga, In) in the 3C-Si host lattice induces a larger contraction compared to the neutral case, presumably because the lattice has fewer degrees of freedom to adjust to the extra electron configuration, resulting in a more significant contraction. The analysis of donors shows that, when positive charges are added, bonds tend to contract as well, but without any clear general trend (see Table 3).
@@
Compared to the 3C-Si case, the 2H-Si host lattice presents a lower symmetry due to the anisotropy along the [0001] hexagonal direction ($C_{3v}$ symmetry). In the pristine structure, for a given Si atom, three equal length first-neighbor bonds in the c-(0001) plane can be defined. The fourth bond is longer and perpendicular to the plane, breaking the tetrahedral symmetry. This arrangement of the atoms increases the degrees of freedom for the structural relaxation around the defect. As shown in the first block of Table 2, the local symmetry of the B0 impurity is $C_{3v}$: there exist three equivalent bonds in the plane and a longer one along the [0001] crystal direction. This is a consequence of three valence electrons of the acceptor trying to keep the three-fold coordination. The difference between the average of the three in-plane bonds and the out-of-plane one is larger than that of the 2H pristine case. As for B-, we observe the same symmetry but slightly more pronounced: an equivalence of the three bonds in the plane and a slight expansion of the perpendicular one with respect to the neutral case. Looking at the second block of Table 2, we can recognize a similar behavior in the case of Al: both Al0 and Al- keep a clear $C_{3v}$ symmetry with three equal length bonds in the plane and a longer one along the $c$ hexagonal direction. As for Ga0 and In0 (third and fourth blocks of Table 2), though it is not very clear, it seems that the local symmetry is closer to the $T_{d}$ cubic one. When a charge is added, this tendency is even more pronounced, especially for In (fourth block of Table 2). On the other hand, donors in 2H-Si induce similar contraction and expansion of the local structure, but with a different impurity symmetry. As reported in the first block of the Table 3, P0 has a $C_{3v}$ symmetry with a difference between the in-plane and the out-of-the-plane bond lengths larger than those of the pure 2H-Si. When a charge is added, the out-of-plane bond length is slightly reduced, while all the in-plane bonds maintain the same value. There is almost no change in the symmetry between neutral and charged cases, keeping a pronounced $C_{3v}$ symmetry. The same behavior can be recognized also for As and Sb (second and third block of Table 3). It is worth to note that the difference between the in-plane and the out-of-plane bond length average is strictly related to the impurity ionic radius being lower for Sb and larger for P.

### v2

@@ Abstract
Recent advances in the characterization of hexagonal-diamond silicon (2H-Si) have shown that this material possesses remarkably different structural, electronic and optical properties as compared to the common cubic-diamond (3C) polytype. Interestingly, despite the wide range of physical properties analyzed, to date no study has investigated impurity energy levels in 2H-Si. Here, we present results of ab initio DFT simulations to describe the effect of p- and n-type substitutional doping on the structural and electronic properties of hexagonal-diamond Si (2H-Si). We first provide a detailed analysis of how a given impurity can assume a different local symmetry depending on the host crystal phase. Then, by studying neutral and charged dopants, we carefully estimate donors and acceptors transition energy levels in 2H-Si and compare them with the cubic-diamond (3C) case. In the case of acceptors, the formation energy is always lower in 2H-Si and is associated with a shallower charge transition level with respect to 3C-Si. On the other hand, donors prefer the cubic phase and have transition energies smaller with respect to 2H-Si. Finally, by employing a simple model based on the 2H/3C band offset diagram, we prove the physical validity of our findings and we show how holes can be used to stabilize the 2H-Si phase. Overall, the described doping properties represent a robust starting point for further theoretical and experimental investigations.
@@ I Introduction
Recent developments and advances in semiconductor growth methods demonstrated that the hexagonal-diamond polytype of Si, Ge and SiGe alloys (2H in the Rasmdell notation [1]) can be fabricated in the form of nanowires [2, 3, 4, 5, 6]. This crystal structure presents a $C_{3v}$ symmetry (P63/mmc space group) with three equal length bonds in the c-plane (0001) and a larger one along the [0001] direction of the hexagonal crystal cell. Compared to the 3C cubic-diamond phase (Fd-3m group symmetry) where four equal length bonds induce a $T_{d}$ symmetry, in the 2H phase the atomic ordering along the $\langle 111\rangle$ cubic direction changes from an ABC to an ABAB stacking.
@@
The main reason for studying these novel group IV polytypes lies in the possibility to observe unexplored electronic and optical properties with respect to those of their 3C counterparts. To give an example, Ge, which in its cubic form is an indirect semiconductor, has a direct band gap in its hexagonal polytype [7, 8]. Yet, hexagonal SiGe alloys [9] and quantum wells [10] have recently attracted a great attention because of their direct band gap and the experimentally observed light emission that can be tailored via the Ge concentration and quantum confinement effects [3, 11, 10]. In the case of Si, while the cubic phase has an indirect band gap of 1.12 eV with a conduction band minimum close to the X point, the 2H structure presents an indirect band gap of about 0.95 eV and a conduction band minimum at the L point [12, 13]. The direct band gap is also reduced from 3.40 eV to 1.63 eV [12]. Indeed, 2H-Si presents a larger optical absorption spectra in the visible range than 3C-Si [14], making it a good candidate for solar panel materials. Furthermore, 3C/2H Si junctions are also of interest for photovoltaics due to their intrinsic tendency to promote charge carrier separation [14, 8, 15].
@@
It is well known that most technologies based on carrier transport in semiconductors (e.g. electronics, embedded systems, light emitting devices and solar cells) become functional only if the material can be appropriately doped. Impurities play a crucial role as they allow to modulate the carrier concentration around the Fermi level. Furthermore, in different type of semiconductor compounds, doping can be employed to stabilize one polytype with respect to the others [16, 17, 18, 19]. For instance, Cai et al. [20] recently demonstrated that the direct band gap 2H-Ge phase can be stabilized by incorporating electrons into the 3C-Ge crystal. Finally, most of the experimental grown nanowires can contain significant concentration of extrinsic dopants due to the substrate atomic contamination. This is, for instance, the case of 2H-Ge and 2H-SiGe alloys [3] fabricated with the crystal transfer methods which usually present a high concentration of As atoms ( $\sim$ 1018 cm-3).
@@
Although the behaviour of impurities in 3C-Si has been widely discussed, to date no study has investigated impurity energy levels in 2H-Si. The experimental characterization of dopants in this polytype still remains challenging because of the difficulties in its fabrications methods. Ab-initio theoretical approaches can be hence an efficient way to investigate this question. A few studies focused on the analysis of the stability of single-point and neutral substitutional defects in 2H-Si [21, 22, 23] but they were essentially limited to the structural stability of defects. Here, we present a comprehensive study carried out via ab initio DFT simulations on the effect of p- and n-type substitutional defects on the structural and electronic properties of 2H-Si. We first provide a detailed analysis on how a given impurity can assume a different local symmetry depending on the host crystal phase. Then, by studying neutral and charged dopants, we carefully estimate donors and acceptors thermodynamic transition levels in 2H-Si and compare them with the 3C-Si case. Finally, by employing a simple model based on the 2H/3C band offset diagram, we show how the carrier concentration can be used to stabilize one phase with respect to the other.
@@ II Methodology
All simulations were performed by spin-polarized ab initio DFT as implemented in the SIESTA code [24], using the generalized gradient approximation (GGA) with the Perdew, Burke, Ernzerhof (PBE) functional [25]. Since the diameter of the 2H nanowires grown experimentally is usually larger than 100 nm, we considered bulk supercells as an accurate and reliable approximation for such systems. Indeed at such large diameters, quantum confinement effect on the band gap can be neglected [26, 27, 28]. Norm-conserving pseudopotentials of Trouillier-Martins type were chosen to replace the core electrons. Valence electrons were instead represented using a double-$\zeta$ (DZ) polarized basis set for Si, and double-$\zeta$ plus two polarization orbitals (DZP) basis sets for impurities. We considered p-type (B, Al, Ga, In) and n-type (P, Sb, As) as common substitutional dopants in 2H-Si. Lattice parameters and band gap of pristine bulk 3C- and 2H-Si were calculated by sampling the Brillouin zone using a 8$\times$8$\times$8 Monkhorst-Pack k-points grid [29]. Table 1 presents the converged results of the lattice parameter and the band gap calculations for both the phases as well as their experimental values. As expected, when compared to experiments, it is clear that both quantities are affected by the error associated to the GGA exchange-correlation functional. We used a denser grid of 25$\times$25$\times$25 k-points to calculate the dielectric constant of both phases. This quantity was needed for the charge correction scheme described below [30]. The values of the in-plane and out-of-plane components, reported in Table 1, are in perfect agreement with other DFT calculations [31, 32] and only slightly overestimate the experimental values [33]. For the doped systems, a supercell of 512 (a 4$\times$4$\times$4 replica of the 8-atoms conventional supercell with orthogonal cell vectors) and of 500 atoms (a 5$\times$5$\times$5 replica of the primitive cell) were used for the 3C and 2H phases, respectively. This choice has guaranteed the convergence of the formation energy value for both neutral and charged impurities as it eliminated the spurious interaction between defect periodical images (see Supplemental Materials (SM) [34]). Such doped supercells correspond to a dopant concentration between 1019 cm-3 and 1020 cm-3, a value that is one order of magnitude larger with respect to the impurity concentration usually measured in 2H-samples [3]. A 2$\times$2$\times$2 k-grid and a real space grid cut-off of 500 Ry were considered to ensure the proper convergence of the results. Self-consistency field cycles were converged with a tolerance of $10^{-5}$ eV. Atomic positions were optimized until the forces were less than $10^{-2}$ eV/Å and the stress lower than $10^{-1}$ GPa. Relaxation was performed adopting a conjugate gradient algorithm that calculates DFT forces through the Hellman-Feynman theorem. Charged structures were relaxed starting from the neutral defect optimized structure. A compensating jellium background was added to avoid divergences due to the defect-defect long-range electrostatic interaction between charged periodic cells.
@@
For all the considered impurities, we calculated their formation energy, $E_{form}$, in a given crystal phase (2H or 3C). This quantity is defined as the energy required to replace a host atom with a neutral or charged defect. In the Zhang and Northrup formalism [35], it reads as follows:
@@
where $E_{total}^{D}$ is the total energy of the doped system, $E_{total}^{prist}$ is the total energy of the pristine system, $\mu_{0}$ is the chemical potential of Si in a given phase while $\mu_{D}$ is that of the defect. $\mu_{0}$ is here defined as the total ground energy per atom of a pristine Si cell whereas $\mu_{D}$ is calculated as the total energy of the isolated impurity atom. However, it is worth to note that there is no univocal criteria to calculate the chemical potential of an impurity as this quantity strongly depends on the synthesis conditions. As a consequence, the calculated value of the formation energy can assume a wide range of values depending on the way $\mu_{D}$ is defined. Nevertheless, since we are interested in the relative stability of a dopant among the two phases, the chemical potential $\mu_{D}$ erases if we refer to the difference between formation energies, $\Delta E^{2H-3C}_{form}=E_{form}^{2H}-E_{form}^{3C}$ rather than in $E_{form}$ value in itself (this also holds because the difference of $\mu_{0}$ among the two phases can be neglected). Yet, referring back to Eq. (1), the term $q(\epsilon_{VBM}+\mu_{e})$ is the charge multiplied by the electron chemical potential, $\mu_{e}$, using as energy reference the valence band maximum, $\epsilon_{VBM}$. $\mu_{e}$ varies from 0 to the electronic band gap. Finally, the term $E_{corr}$ corresponds to the Freysoldt-Neugebauer-Van de Walle’s finite-size correction [30] that cancels artificial interactions between charged defects and that has been proven to give very accurate results for different type of defects [36]. As other proposed schemes [37, 38, 36, 39], this approach implies the solution of the Poisson equation for a model system and the alignment of the electrostatic potentials to derive the formation energy correction. In particular, solving the Poisson equation for the periodic model potential needs the calculation of the dielectric tensor profile of the material [39], whose static components are presented in Table 1.
@@
For each dopant the formation energy plots are reported for different values of $q$ (-1, 0 and +1) and they are linear functions of the chemical potential as shown in Eq. 1. The intersection of the formation energy lines for two different charge states $q_{1}$ and $q_{2}$ is the thermodynamic transition level calculated through Eq. 2
Transition levels $\epsilon(q_{1}/q_{2})$ correspond to thermal ionization energies that can be measured by deep level transient spectroscopy (DLTS) [40]. We employed Eq. (2) to calculate both acceptors, $\epsilon_{A}(q_{1}/q_{2})$ and donor transition levels $\epsilon_{D}(q_{1}/q_{2})$. It is worth to note that transition levels are not affected by the dopant chemical potential which assumes the same value for the charged and neutral impurity and hence cancels out in Eq. (2).
In the following analysis, Eq. (2), which depends on the valence band maximum of each phase taken separately, allows to compare the doping with a given impurity of the 2H and the 3C phase of Si. On the other hand, to give an accurate quantitative estimation of transition levels, in particular in the case of the phase coexistence [4, 5, 14], we align $\epsilon$($q_{1}$/$q_{2}$) to a common reference level, i.e. the vacuum level $\Phi$ (see SM for the details of the calculations [34]). Indeed, many works have demonstrated that, even if within a semi-local approach, energy levels of dopants generally remain preserved if they are referred to a common well defined potential [41, 42, 43]. This choice improves the quantitative comparison among the two phases. On the basis of this consideration, Eq. (2) can be modified as:
@@ III Structural properties
We start our discussion by a detailed structural analysis of the relaxed doped structures. Table 2 and 3 report the first-neighbors bond lengths of group III and group V impurities in the cubic and hexagonal crystals. In the 3C-Si, both neutral donors and acceptors maintain the tetrahedral symmetry ($T_{d}$) forming four first-neighbor bonds of equal length. Variations in bond lengths compared with the pristine system are uniform and depend on the mismatch between the covalent radius of the impurity and that of the host element, as already reported by Amato et al. [21, 22]. The bond contractions and expansions discussed here agree with previous results. As for smaller defects, such as B and P, the first neighbors impurity bonds are compressed in respect to those of the pristine cell (see first block of Table 2 and  3) while for all the other reported defects (Al, Ga, In, Sb and As in Table 2 and  3) they tend to expand since their covalent radius is larger than those of Si. More in general, we can say that impurities in the 3C-Si phase induce an expansion (contraction) of their first neighbors bonds and, consequently, there is a noticeable contraction (expansion) between the first and second neighbors atomic layers that gradually fades within the crystal. The charged impurities also preserve the cubic $T_{d}$ symmetry. As for B-, the extra electron does not induce any appreciable change in the bond distances (see first block of Table 2). Al-, Ga-, and As- first neighbors distances are only slightly reduced compared to the neutral case (see Table 2). These tiny bond length variations are proportional to the impurity size, being more pronounced for larger dopants. Therefore, adding an extra electron to large radius acceptors (Al, Ga, In) in the 3C-Si host lattice induces a larger contraction compared to the neutral case, presumably because the lattice has fewer degrees of freedom to adjust to the extra electron configuration, resulting in a more significant contraction. The analysis of donors shows that, when positive charges are added, bonds tend to contract as well, but without any clear general trend (see Table 3).
@@
Compared to the 3C-Si case, the 2H-Si host lattice presents a lower symmetry due to the anisotropy along the [0001] hexagonal direction ($C_{3v}$ symmetry). In the pristine structure, for a given Si atom, three equal length first-neighbor bonds in the c-(0001) plane can be defined. The fourth bond is longer and perpendicular to the plane, breaking the tetrahedral symmetry. This arrangement of the atoms increases the degrees of freedom for the structural relaxation around the defect. As shown in the first block of Table 2, the local symmetry of the B0 impurity is $C_{3v}$: there exist three equivalent bonds in the plane and a longer one along the [0001] crystal direction. This is a consequence of three valence electrons of the acceptor trying to keep the three-fold coordination. The difference between the average of the three in-plane bonds and the out-of-plane one is larger than that of the 2H pristine case. As for B-, we observe the same symmetry but slightly more pronounced: an equivalence of the three bonds in the plane and a slight expansion of the perpendicular one with respect to the neutral case. Looking at the second block of Table 2, we can recognize a similar behavior in the case of Al: both Al0 and Al- keep a clear $C_{3v}$ symmetry with three equal length bonds in the plane and a longer one along the $c$ hexagonal direction. As for Ga0 and In0 (third and fourth blocks of Table 2), though it is not very clear, it seems that the local symmetry is closer to the $T_{d}$ cubic one. Two bonds in the plane are larger than the third one, and even larger than the perpendicular interatomic distance, with this effect being more pronounced for In due to its larger atomic size. When a charge is added, the tendency toward $T_{d}$ symmetry becomes more evident, especially for In (fourth block of Table 2). Possible factors that could explain this deviation from the expected bond length pattern include local distortions, impurity size mismatch as well as electronic wave function symmetry breaking. On the other hand, donors in 2H-Si induce similar contraction and expansion of the local structure, but with a different impurity symmetry. As reported in the first block of the Table 3, P0 has a $C_{3v}$ symmetry with a difference between the in-plane and the out-of-the-plane bond lengths larger than those of the pure 2H-Si. When a charge is added, the out-of-plane bond length is slightly reduced, while all the in-plane bonds maintain the same value. There is almost no change in the symmetry between neutral and charged cases, keeping a pronounced $C_{3v}$ symmetry. The same behavior can be recognized also for As and Sb (second and third block of Table 3). It is worth to note that the difference between the in-plane and the out-of-plane bond length average is strictly related to the impurity ionic radius being lower for Sb and larger for P.